Sandy pulls a boat a distance of 15.8 meters along a dock using a rope that makes an angle of 20.0o with the horizontal. If she uses a force of 888 N, calculate work?
1
Expert's answer
2020-11-20T07:14:22-0500

By the definition of the work done, we have:


"W=Fscos\theta=888\ N\cdot 15.8\ m\cdot cos20.0^{\circ}=13184\ J."

Answer:

"W=13184\ J."